OpenStudy (anonymous):

I don't understand the example from textbook --------------------------- Starting with basic {1,x,x^2} , we can use Gram-Schmidt process to generate an orthnormal basis. ||1||^2=<1,1>=3 why is it 3?

OpenStudy (beginnersmind):

Either a typo or they are using a very unusual norm function. Maybe try reading back a few pages?

OpenStudy (phi):

What is their definition of <1,1>? I would expect it to = 2 using an integral over the interval -1 to 1

OpenStudy (anonymous):

more about problem Find orthonormal basics for P_3 if the inner product on P_3 is defined by \[\sum _{i=1}^3 p\left(x_i\right)q\left(x_i\right)\] where x1=-1 , x2=0, x3=1.

OpenStudy (anonymous):

@Zarkon , help

OpenStudy (dumbcow):

sorry, not as familiar with linear algebra and working with orthonormal basis in other words i don't remember and would have to study up on it

OpenStudy (dumbcow):

@myininaya , could you help

OpenStudy (anonymous):

@Thomas9 , can you help?

OpenStudy (phi):

I think(?) p and q are one of 1, x, x^2 they define the inner product for 1 as 1*1+1*1+1*1 = 3 (i.e. 1 is not a function) for p=1 q=x we get 1*(-1) + 1*0+ 1*1 =0 for p=x , q=x we get -1*-1 +0*0+1*1= 2

OpenStudy (phi):

with this definition you can apply Gram-Schmidt

OpenStudy (phi):

I think you get b1= 1/sqrt(3) b2= x/sqrt(2) now you have to find b3

OpenStudy (anonymous):

I don't understand how that innerproduct works: You have two vectors p and q, and you multiply element wise and take the sum? Just like the regular innerproduct?

OpenStudy (anonymous):

If that's the case, isn't {1,x,x^2} already an orthonormal basis?

OpenStudy (phi):

p(x) is a polynomial that you evaluate at -1, 0, 1

OpenStudy (anonymous):

I see, in that case you're correct so far.

OpenStudy (anonymous):

So ||1||^2= (1,1) = 1*1+1*1+1*1=3 ||1||=sqrt(3) So first step is to divide the the first vector 1 by sqrt(3). ((1,1) is the innerproduct, not sure if that notation is common)

OpenStudy (anonymous):

x is already orthogonal to 1, so that's easy. ||x||=sqrt(2), so you have to divide that vector by sqrt(2)

OpenStudy (anonymous):

Thanks, I have much better understading now

OpenStudy (anonymous):

Thanks, Phi and Thomas

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

to finish this, x^2 is orhogonal to x, but not to 1, so you have to add something to that vector and then divide by its length.

OpenStudy (phi):

that last bit looks ugly

OpenStudy (anonymous):

isn't it x^2-(x^2,1)1/|| x^2-(x^2,1)1|| I'm not sure haven't used those formulas for a while.

OpenStudy (phi):

yes. but evaluate it

OpenStudy (anonymous):

(x^2-2)/sqrt(6)?

OpenStudy (phi):

tho the 1's should be unit length x^2 dot 1/|1| time 1/|1| (if that makes sense)

OpenStudy (anonymous):

I see

OpenStudy (phi):

yes, that is good. I think that has unit length.

OpenStudy (phi):

but it is not orthogonal to 1 is it?

OpenStudy (anonymous):

I don't think so no.

OpenStudy (anonymous):

if a=1/sqrt(3) Then the third vector should be x^2-(x^2,a)a/|| x^2-(x^2,a)a||

OpenStudy (anonymous):

but that's the same as x^2-(x^2,1)/3/|| x^2-(x^2,1)/3||

OpenStudy (anonymous):

which equals (x^2-2/3)/sqrt(2/3) I think.

OpenStudy (anonymous):

It's orthogonal to both 1 and x and has lenght 1, so that should be it.

OpenStudy (anonymous):

i didnt read thru everyone elses post but do u want me to solve this question?